What is the probability of rolling a sum less than 5 or greater than 10 when tossing two dice?
Answer by rajagopalan(174)

p less than 5
1+3
2+2
3+1
1+2
2+1
1+1
6 occurances in 36 possible outcomes=6/36
..
p greater than 10
5+6
6+5
6+6
3 occurances in 36 possible outcomes=3/36
..
so for either <5 or >10 add the probabilities
6/36+3/36=9/36=1/4
answer=1/4